### v1

# Domain-Aware Tensor Network Structure Search

Abstract
Tensor networks (TNs) provide efficient representations of high-dimensional data, yet identification of the optimal TN structures, the so called tensor network structure search (TN-SS) problem, remains a challenge. Current state-of-the-art (SOTA) algorithms are computationally expensive as they require extensive function evaluations, which is prohibitive for real-world applications. In addition, existing methods ignore valuable domain information inherent in real-world tensor data and lack transparency in their identified TN structures. To this end, we propose a novel TN-SS framework, termed the tnLLM, which incorporates domain information about the data and harnesses the reasoning capabilities of large language models (LLMs) to directly predict suitable TN structures. The proposed framework involves a domain-aware prompting pipeline which instructs the LLM to infer suitable TN structures based on the real-world relationships between tensor modes. In this way, our approach is capable of not only iteratively optimizing the objective function, but also generating domain-aware explanations for the identified structures. Experimental results demonstrate that tnLLM achieves comparable TN-SS objective function values with much fewer function evaluations compared to SOTA algorithms. Furthermore, we demonstrate that the LLM-enabled domain information can be used to find good initializations in the search space for sampling-based SOTA methods to accelerate their convergence while preserving theoretical performance guarantees.
1 Introduction
The exponential increase in the volume and richness of available data has led to the widespread use of multi-way arrays, often represented as higher-order tensors. Tensor network decomposition methods aim to represent higher-order tensors in “super-compressed” formats through smaller-sized components, by effectively capturing cross-modal latent patterns and correlations. These methods have been applied across various fields, including machine learning [1, 2, 3], signal processing [4], computer vision [5, 6], and quantum physics [7, 8]. The success of tensor network decomposition techniques is closely linked to their ability to mitigate the “curse of dimensionality”, which is achieved by decomposing higher-order data into lower-order factors.
However, tensor network (TN) practitioners face significant challenges related to choosing the most appropriate TN structure, the so called tensor network structure search (TN-SS) problem, which has been proven to be NP-hard [9, 10]. TN-SS involves determining the optimal TN parameters, such as TN ranks, TN topology, and TN mode permutations [11].
Existing methods to efficiently resolve the TN-SS problem (see Section 1.1) include Bayesian inference [12], spectrum methods [13], reinforcement learning [14], and discrete [11, 10] and continuous optimization [15], along with greedy search and random sampling. To date, sampling-based methods [16, 11, 10, 17], whose workflows are illustrated in Figure 1a and 1b, have demonstrated the best performance in addressing the TN-SS problem.
However, these methods require large number of evaluations (of the training and testing data), are prone to getting stuck in local minima [10], and lack transparency in their found structure-related parameters. Critically, a large number of evaluations needed to optimize the objective function leads to a high computational cost. We hypothesize that these weaknesses of the current methods are due to the failure to utilize the rich domain information inherent in real-world tensors, such as the mode information. To this end, we ask ourselves:
Can we utilize the intrinsic domain information in tensor data to significantly reduce the number of evaluations required to solve the TN-SS problem, while providing domain-aware explanations for the identified TN solutions?
Figure 1: Comparison of SOTA tensor network structure search (TN-SS) methods. (a) TN-SS algorithms with theoretical guarantees: TNLS [11] & TnALE [10]. (b) tnGPS [17], a prompting pipeline which uses LLMs to generate new TN-SS algorithms. (c) tnLLM (ours), which uses domain information about the tensor data and LLM reasoning to solve the TN-SS problem and generate explanations for the identified TN structure.
To address this question, we propose a domain-aware large language model (LLM)-guided TN-SS framework, termed tnLLM. Within this framework, an LLM is utilized to initialize the TN structure based on domain information about the relationships between tensor modes. Then, the reasoning capabilities of the LLM are used to navigate the search space effectively, in order to achieve good optimization of the objective function with very few evaluations.
Our proposed framework is found to achieve significant speed-ups over current state-of-the-art (SOTA) methods in terms of the number of evaluations, due to its ability to find good TN structure initializations through the use of domain information. Moreover, tnLLM generates practically meaningful explanations for the TN solutions, thus offering transparency in the identified structures. This is particularly beneficial for tensor practitioners who lack deep expertise in a specific data domain, as it enables them to both comprehend the interactions between tensor modes and trust the discovered TN structures. It also allows the identified structures to be verified by domain experts.
[…]
  • We propose tnLLM, a novel domain-aware LLM-guided TN-SS framework, which achieves performance on par with SOTA methods while requiring much fewer evaluations;
  • To the best of our knowledge, our framework is the first to utilize domain information inherent in real-world tensor data to address the TN-SS problem. This enables the generation of domain-aware explanations that allows practitioners to verify the identified TN structures.
Figure 2: The workflow of the proposed tnLLM framework consists of two key stages commonly adopted by SOTA TN-SS methods [11, 10, 17]: TN-initialization and TN-discovery. The proposed tnLLM efficiently utilises the domain information in tensor data and uses an LLM to guide these two stages.
1.1 Related work
Tensor network structure search (TN-SS).
Compared to traditional tensor decompositions [18, 19, 20, 21, 22], which have pre-defined tensor network structures, the TN-SS problem focuses on finding custom tensor networks, which have been shown to achieve higher parameter efficiency and are an important paradigm of tensor decompositions [10, 11, 16, 17]. Various approaches have been proposed to address the TN-SS problem, including Bayesian inference [12], spectrum methods [13], reinforcement learning [14], program synthesis [23], and continuous optimization [15]. Sampling-based methods [10, 11, 16, 17], which fall under discrete optimization methods, have demonstrated superior performance compared to other approaches to address the TN-SS problem. Among these, TNLS [11] and TnALE [10] achieve state-of-the-art (SOTA) performance, with TnALE accelerating TNLS by slightly reducing the number of evaluations required to optimize the objective function. While both TNLS and TnALE adopt a “local-search” scheme, this acceleration is achieved by the alternating variable update framework proposed in TnALE. However, these methods still face challenges, such as a large numbers of function evaluations required to converge, difficulty in balancing exploration and exploitation, and a lack of transparency in the found TN structures.
More recently, tnGPS [17] has emerged as an approach that uses LLMs to generate sampling-based TN-SS algorithms, demonstrating performance comparable to SOTA methods. Despite being closest to our work, tnGPS focuses on generating TN-SS algorithms based on sampling-based heuristics and does not incorporate any domain information in the tensor data. As such, tnGPS suffers from the same limitations as existing sampling-based methods, including a lack of explainability in the identified structure-related parameters.
To this end, our work proposes to harness the domain information in tensor data to solve the TN-SS problem. We achieve this by using the domain knowledge and the inherent reasoning capabilities of LLMs to directly infer tensor network structures. Facilitated by our carefully designed prompting pipeline (see Section 3), the proposed framework is shown not only to require significantly fewer evaluations, but also to produce domain-aware and verifiable solutions to the TN-SS problem.
Reasoning with large language models.
The rise of transformer-based LLMs, pretrained on vast text corpora, has demonstrated a remarkable capacity for “reasoning” [24]. This reasoning ability is further enhanced when LLMs are guided by task-specific prompting strategies, such as chain-of thought [25, 26, 27]. These strategies enable LLMs to generate text effectively for tasks such as arithmetic reasoning and factual knowledge retrieval, leading to exceptional performance in complex question-answering and tasks requiring real-world knowledge [28, 29]. These findings suggest that, through pretraining on diverse textual data, LLMs encode rich knowledge about real-world relationships, which they can effectively leverage to perform various downstream tasks [30]. In this work, we utilize the real-world domain knowledge and “reasoning” capabilities of LLMs to directly infer suitable TN structures.
2 Preliminaries
In this section, we first summarize the notations for TNs and review the TN-SS problem. Then, we provide a brief overview of SOTA sampling-based methods for TN-SS such as TNLS [11], TnALE [10], and tnGPS [17]. Their workflow guides the design of our proposed framework.
2.1 Tensor network structure search (TN-SS)
We first provide the definition of TN-SS through its application in higher-order data decomposition. An optimal solution to the TN-SS problem aims to find the best trade-off between identifying the most compressed TN structure while preserving the expressivity of the TN [31]. Let $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ be a non-zero order-$N$ tensor, with its Frobenius norm as $\|\mathcal{X}\|_{F}$. Tensor networks can be represented using the graphical notation [32, 33], whereby a tensor network $G$ is represented using a set of $N$ vertices, $V$, and a set of edges, $E$, i.e., $G=(V,E)$. Each vertex represents a decomposed core tensor, and the closed edges between two core tensors are generalized higher-order matrix multiplications, termed tensor contractions [2]. Closed edges have assigned TN-ranks, $\mathbf{r}\in\mathbb{Z}_{+}^{E}$, which indicate the degree of connections between different pairs of connected vertices. Therefore, the properties of a TN structure can be fully expressed by the combination $(G,\mathbf{r})$.
Similar to SOTA TN-SS methods[16, 10, 11, 17], the discrete optimization problem of TN-SS is formalized as a minimization of the objective function, which is a linear sum of the complexity of the TN structure (e.g., compression rate) and the TN expressivity (e.g., approximation error), and is given by
$\min_{(G,\mathbf{r})}\ \ln\Bigg{(}\phi(G,\mathbf{r})\\ +\frac{\lambda}{L}\min_{\{\mathcal{V}_{l,i}\}_{i=1}^{N}}\sum_{l=1}^{L}\frac{\left\|\mathcal{X}_{l}-TNC\left(\{\mathcal{V}_{l,i}\}_{i=1}^{N};(G,\mathbf{r})\right)\right\|_{F}}{\left\|\mathcal{X}_{l}\right\|_{F}}\Bigg{)}$ (1)
where $L$ represents the number of tensor samples in the dataset, and the first term $\phi(G,\mathbf{r})$ measures the TN structure complexity. The second term in Equation (1) measures the expressivity of the TN through the relative squared error (RSE) between the original tensors $\{\mathcal{X}_{l}\}_{l=1}^{L}$ and their TN approximations $\left\{TNC\left(\{\mathcal{V}_{l,i}\}_{i=1}^{N};(G,\mathbf{r})\right)\right\}_{l=1}^{L}$, where $TNC(\cdot)$ stands for the tensor contraction operation of the entire tensor network. The pair $(G,\mathbf{r})$ characterizes how the vertices, $\{\mathcal{V}_{l,i}\}_{i=1}^{N}$, are contracted together to approximate the original tensor $\mathcal{X}_{l}$. The coefficient $\lambda$ is a positive non-zero scaling factor which balances the trade-off between the model complexity and model expressivity. Note that [16, 33] pointed out that the TN-SS problem is conveniently equivalent to the TN rank search problem of a fully connected TN [34].
Algorithm 1 : Sampling-based Algorithms for TN-SS
1:Initialize:
2:$N_{Iter}$triangleright Maximum number of iterations
3:$N_{Sample}$triangleright Number of samples per iteration
4:$P\leftarrow[]$triangleright Historical TN structures
5:$H\leftarrow[]$triangleright Discovered TN structures in each iter.
$\mathcal{F}(\cdot)\leftarrow$ Equation (1) $\triangleright$ Objective function
7:$(G,\mathbf{r})=\textit{Initial TN Structure}$triangleright TN-initialization
8:
9:Algorithm:
10:for $n=1$ to $N_{Iter}$ do
11:    $H\leftarrow$ $N_{Sample}$ TN structures sampled in the neighborhood of $(G,\mathbf{r})$. $\triangleright$ TN-discovery
12:    $P\leftarrow P\cup H$
13:    if  $\exists(\hat{G},\hat{\mathbf{r}})\in P$ such that $\mathcal{F}(\hat{G},\hat{\mathbf{r}})<\mathcal{F}(G,\mathbf{r})$ then
14:         $(G,\mathbf{r})\leftarrow(\hat{G},\hat{\mathbf{r}})$.
15:    end if
16:    if Converged then
17:         return $(G,\mathbf{r})$
18:    end if
19:end for
20:Output: $(G,\mathbf{r})$
2.2 SOTA sampling-based algorithms for TN-SS
Algorithm 1 summarizes the current SOTA algorithms for TN-SS such as [11, 10, 17]. They follow a three-step search process: TN-initialization $\rightarrow$ TN-discovery via sampling in the search neighborhood $\rightarrow$ Updating the center of the search neighborhood. Existing methods ignore the inherent domain information in real-world tensor data, which calls for the development of a framework that can effectively utilize domain knowledge in TN-initialization and TN-discovery to improve performance.
To this end, our proposed framework introduces domain-aware LLM-guided TN-initialization and TN-discovery. By doing so, it addresses the limitations of existing methods, such as high computational costs caused by the large number of evaluations ($N_{Iter}\times N_{Sample}$) required and the tendency to get stuck in local minima due to difficulties in balancing the exploration-exploitation trade-off. Importantly, by incorporating domain information, our tnLLM framework provides practically meaningful explanations for the identified TN structures, a feature absent in current approaches.
3 tnLLM: a domain-aware framework for solving TN-SS
This section presents tnLLM, a domain-aware LLM-guided framework that efficiently solves the TN-SS problem with very few evaluations and verifiable solution explanations. We detail the role of each component in the prompting pipeline and the guidelines followed by each prompt.
Workflow of tnLLM.
Figure 2 illustrates the workflow of the proposed tnLLM framework. The “Behavior-directive” prompt is used to supply the problem specification and guide the general-purpose LLM into a model tailored for solving the TN-SS problem. Inspired by current sampling-based SOTA algorithms, this domain-informed LLM is then employed for LLM-guided TN-initialization and TN-exploration.
During the TN-initialization stage, the “Task-directive” prompt is used to propose a strong initial TN structure by using the provided domain information. In the TN-discovery stage, the last and best-identified TN structures are used in conjunction with the “Optimization-directive” prompt to guide the LLM in navigating the search space effectively. This iterative process leads to the refinement of the TN structure and yields improved objective function values over successive evaluations.
Prompting pipeline.
The interactions between the carefully designed user prompts and the LLM assistant follow an automated structured dialogue [35, 36]. This dialogue-based approach is intuitive for generating stepwise conversational reasoning and particularly useful for iterative tasks such as the TN-SS problem. To further improve the ability of the LLM to perform complex reasoning, we employ the chain-of-thought prompting [25]. This prompting strategy breaks down complex tasks into a series of intermediate reasoning steps and thus enables the model to solve the problem by addressing smaller, simpler sub-tasks sequentially. In the proposed tnLLM framework, three distinctly purposed prompts are employed, each serving a unique role. The full prompts are given in Appendix B.
“Behavior-directive” prompt.
This is the system message in the dialogue. Its primary objective is to establish the role of the LLM as an expert in the TN-SS problem (orange section of Figure 3(a)) and guide its behavior throughout the dialogue (green section of Figure 3(a)). Additionally, it sets the context by outlining the TN-SS problem and the objective function (blue section of Figure 3(a)).
“Task-directive” prompt.
This is the first non-system prompt supplied to the LLM and serves as the basis for LLM-guided TN-initialization. It is designed to thoroughly explain the domain information about the tensor structure, including details such as the number of modes and the specific information about each mode (orange section of Figure 3(b)). Furthermore, it guides the model’s thought process and ensures that its responses align with the task requirements by instructing it to reason step-by-step and utilize domain knowledge (blue section of Figure 3(b)).
(a) “Behavior-directive” prompt
(b) “Task-directive” prompt
Figure 3: Structure and components of (a) the “Behavior-directive” prompt, which frames the LLM’s role as a domain expert, and (b) the “Task-directive” prompt for LLM-guided TN-initialization.
“Optimization-directive” prompt.
This is the iterative prompt supplied to the LLM in all subsequent evaluations and serves as the basis for LLM-guided TN-discovery. Its purpose is to guide the LLM to efficiently navigate the search space by incorporating context from the best and previous evaluations (orange section of Figure 4). The prompt explains how refinements to the identified TN structure influence the objective function value and instructs the model to leverage step-by-step reasoning and domain knowledge to optimize the objective function (blue section of Figure 4). At the same time, it encourages the LLM to both explore and exploit new solutions to reduce the likelihood of the method being stuck in local minima (purple section of Figure 4).
Figure 4: Structure and components of “Optimization-directive” prompt used for LLM-guided TN-discovery.
Efficient and effective search space navigation.
A critical factor in the performance of tnLLM, both in terms of evaluations and objective function value, is the incorporation of memory (orange section of Figure 4) and the encouragement of search space "exploration and exploitation" (purple section of Figure 4) in the “Optimization-directive” prompt. By including the best and most recent TN solutions, along with their respective objective function values and the domain information about mode interactions, the LLM gains a better understanding of the search space and can make more informed decisions. Additionally, by encouraging the model to explore new solutions when necessary and exploit good solutions already found assists the optimization process, by reducing the possibility of the model getting stuck in local minima, a common caveat of sampling-based methods.
Output format specification.
As observed in Figures 3(b) and 4, in addition to the goal-oriented components of each prompt, an essential feature (highlighted in green) of both the “Task-directive” and “Optimization-directive” prompts is the specification of the exact output format of the TN-SS solution. This includes defining both the sequence in which the LLM should present its response and the precise format of the TN-SS solution. To meet these requirements, the prompt shown in Figure 5 was developed through experimentation and a trial-and-error process with the LLM. This step serves to eliminate “hallucinations” in the LLM output and enables a fully automated prompting pipeline without the need for any human intervention during the entire iterative optimization process.
Figure 5: The prompt used to specify the LLM output format.
Remark 3.1.
Both tnGPS and tnLLM use LLMs to solve the TN-SS problem, however, the two approaches are fundamentally different. tnGPS uses the LLM to generate new sampling-based TN-SS algorithms by reviewing existing algorithms, causing the discovered methods to inherit similar limitations. In contrast, tnLLM uses the LLM to directly infer TN structures to solve the TN-SS problem by incorporating the rich domain information in real-world tensor data. In turn, this allows tnLLM to provide explanations for the TN solutions, thereby adding transparency to the identified TN structures.
4 Experimental results
In this section, the performance of tnLLM was evaluated against the SOTA sampling-based methods in tensor decomposition tasks in terms of the number of evaluations and objective function value. Our results demonstrate that tnLLM achieves comparable objective function values while delivering significant speed-ups. Moreover, we demonstrate its ability to generate domain-consistent explanations for the identified TN structures. Finally, an ablation study was conducted to evaluate the effectiveness of the proposed prompting pipeline and assess the framework’s robustness across different LLMs.
Data preparation.
Given the flexibility of our method to handle tensor data of any order, we evaluated its performance on three types of tensor data with varying sizes and dimensionalities across different domains. In particular, we tested on sets of order-$3$ RGB images data (each tensor sample has size of $144\times 176\times 3$), order-$4$ RGB videos data (each tensor sample has size of $144\times 176\times 3\times 10$), and order-$5$ financial time-series data (each tensor sample has size of $3\times 6\times 3\times 4\times 5$), which, to the best of our knowledge, is the largest dataset in terms of number of samples ever considered in the TN-SS problem. All entries were standardized to values in $[0,1]$, with $80\%$ of each dataset used for training, and the remaining $20\%$ for testing. More details about the data can be found in Appendix C.
Settings of tnLLM.
In Equation (1), we set $\lambda=10$ and used the same compression ratio function, $\phi$, as in previous sampling-based methods [16], defined as the ratio between the number of parameters in the compressed TN format and the original tensor. The maximum number of evaluations was set to $500$ for the images and videos datasets, and $250$ for the financial time-series dataset. An early stopping criterion with a patience of 5 and a delta of 0 was applied. For all experiments, the LLM model GPT-4o (gpt-4o-2024-08-06) [37] was used, with a temperature setting of 0.2.
Implementation details.
In all experiments, we also implemented the three SOTA sampling-based TN-SS algorithms, namely TNLS [11], TnALE [10] and tnGPS [17]. Since the vanilla TNLS is designed to search only for the permutation of a TN, we extended it to fit the settings of TN-SS. For fair comparisons, all baseline methods were evaluated using the same objective function and maximum number of evaluations, with one evaluation defined as a single pass through the entire training/ testing dataset. The full implementation details are provided in Appendix D.
Numerical results.
We ran tnLLM five times and examined the mean and standard deviation for both the objective function values and the number of evaluations to demonstrate its robustness. Observe from Table 1 that tnLLM achieved performance on par with SOTA sampling-based algorithms in both training and test objectives across all data types, as measured by the objective function. Importantly, tnLLM minimizes the objective function with significantly fewer evaluations, requiring up to 78×78$\times$78 × fewer evaluations than TNLS, 41×41$\times$41 × fewer than TnALE, and 110×110$\times$110 × fewer than tnGPS.
Remark 4.1.
We used the number of evaluations, rather than runtime, as our primary metric for computational cost comparisons, as it is a standard hardware-invariant metric adopted by SOTA models [11, 10, 17]. Nevertheless, in tnLLM, LLM inference accounted for only 9.41%, 1.19%, and 2.12% of the total runtime on the images, videos, and time-series datasets, respectively. Given the low runtime required by tnLLM, this translates to less than 1 extra evaluation.
Table 1: Performance comparison across different datasets. The values on the left give the lowest training objective function values and their corresponding testing objective function values. The values in [square brackets] give the number of evaluations to first achieve the best training objective function value. For robustness assessment of tnLLM, we report the average and standard deviation of both the objective function value and the number of evaluations across 5 independent runs. For both metrics, a lower value is better. The best values are denoted in bold. The second best values are underlined.
Data Type TNLS TnALE tnGPS tnLLM (Ours)
Images Train -0.66 [114] -0.65 [81] -0.66 [438] -0.63±$\pm 0.01$ [4.0$\pm$1.9]
Test -0.47 -0.46 -0.44 -0.48$\pm$0.05
Videos Train -1.64 [484] -1.66 [254] -1.65 [175] -1.63±$\pm 0.01$ [6.2$\pm$3.5]
Test -1.72 -1.72 -1.71 -1.70$\pm$0.02
Time-series Train -0.45 [218] -0.47 [177] -0.39 [38] -0.42±$\pm 0.02$ [5.6$\pm$4.4]
Test -0.43 -0.47 -0.40 -0.41$\pm$0.02
Figure 6: Summarized generated explanations of tnLLM for the initialized TN structure, with and without the incorporation of domain information in the prompts.
Domain-aware explanations.
To assess the domain-relevance of the identified TNs, we manually verified the explanations provided for the identified TN structures in all three datasets tested. The explanations of the tnLLM model for the initialized TN structure, with and without domain knowledge, are provided in Figure 6. Observe that without domain information, although the model is inherently capable of reasoning, its lack of domain information about the modes in the real-world tensors causes it to default to random assumptions, likely influenced by its pre-training data, such as assuming a mode as resembling time or frequency. As a result, the explanations it provides are entirely incorrect from a practical perspective and are therefore ineffective in solving TN-SS and guiding TN practitioners.
In contrast, incorporating domain information enables the model to identify the relationships between modes (highlighted in blue in Figure 6) and, more importantly, relate them to real-world principles in vision and finance domains (highlighted in green). The model also shows consistent reasoning in its rank suggestions: when describing a rank as low, medium, or high, it quantitatively selects a value that is coherent relative to other ranks within the same tensor. Moreover, it adjusts these values based on comparisons across different mode pairs, for example by assigning similar ranks to the Height–RGB and Width–RGB modes in the images dataset, and a lower rank to the RGB–Frames mode in the videos dataset. This demonstrates the model’s understanding of the tensor mode relationships based on domain information and supports the validity of the identified TN structure. Consequently, tnLLM offers explanations that are practically useful in helping domain experts to verify and trust the discovered structures. To demonstrate robustness, the complete summarized explanations across three runs for all tested datasets are provided in Appendix E.
4.1 Accelerating SOTA methods with tnLLM
In the first part of Section 4, we have demonstrated that tnLLM achieves comparable objective function values compared to SOTA sampling-based algorithms, while generating domain-aware explanations for the identified TN structures and significantly reducing the required number of function evaluations. However, due to the black-box nature of LLMs, no theoretical analysis can be provided for the evaluation efficiency of tnLLM, in contrast to TNLS [11] and TnALE [10].
At the same time, while sampling-based methods follow a “local-search” scheme within a neighborhood, there is no guarantee on the minimum number of evaluations required to find a “good” neighborhood. Moreover, poor initialization significantly increases the number of evaluations needed. To address this, we construct a hybrid algorithm that first runs tnLLM for 10 evaluations to leverage its domain knowledge and reasoning capabilities in order to identify a strong initialization point in the “global-search” stage. Sampling-based methods then perform “local-search” in the identified neighborhood, combining the speed-up benefits of tnLLM with the theoretical guarantees of sampling-based approaches.
Figure 7: Experimental results across all 3 datasets show that both TNLS and TnALE are significantly accelerated with an initial center of search neighborhood found by tnLLM. We ran TNLS and TnALE with tnLLM initialization in all 3 datasets over 50 evaluations, which include the plotted 10 evaluations performed by tnLLM. Vanilla TNLS and TnALE performed 500 evaluations for both the images and videos datasets and 250 evaluations for the financial time-series dataset if not converged.
Numerical results.
Figure 7 demonstrates the performance difference of TNLS and TnALE with and without a tnLLM-initialized structure. The “global-search” capability of tnLLM significantly sped up the iterative minimization process. During “local-search”, both TNLS and TnALE further improved upon the structures found in “global-search” by tnLLM. Overall, the tnLLM-initialized algorithms achieved nearly identical objective function values with up to 23×23$\times$23 × fewer evaluations compared to vanilla TNLS, and up to 13×13$\times$13 × fewer evaluations compared to vanilla TnALE. The full numerical performance comparison is provided in Appendix F.
4.2 Ablation study
Removal of domain information.
To assess the effectiveness of injecting domain information into the TN-SS problem, we removed the carefully designed structured prompts of the tnLLM framework, equivalent to using the LLM to solve the TN-SS problem without any priors on the domain information. Without domain information, the LLM generates a poorly initialized TN structure that is 80.3% worse in terms of objective function value, due to its lack of information about the different modes, and requires 10.4×10.4$\times$10.4 × more evaluations to converge. Moreover, as the model defaults to random assumptions about the modes, as illustrated in Figure 6, it was found to produce uniformly connected TN structures throughout the minimization process. This restricts the search space to a much smaller set of candidate solutions that are not domain-meaningful from a practical perspective.
Selection of LLM models.
To examine how the choice of LLM affects the performance of tnLLM, we compared our baseline model, GPT-4o (gpt-4o-2024-08-06), against GPT-4.5 (gpt-4.5-preview-2025-02-27), GPT4o-mini (gpt-4o-mini-2024-07-18) and GPT-3.5 (gpt-3.5-turbo-1106). Observe from Table 2 that due to the carefully designed prompts and the structure of the overall framework, tnLLM is robust to the choice of the LLM used. While there are small variations in the achieved objective function values, the overall performance remained consistent across all three datasets in both training and test sets. Furthermore, the number of evaluations required by all LLMs falls within the mean ± standard deviation range reported in Table 1. It is important to note, however, that weaker models, and in particular GPT-3.5, are more prone to “hallucinations”. For example, we observed instances where the model misinterpreted the optimization objective, mistakenly assuming that higher objective values were always preferable.
Table 2: Ablation study of using different LLMs within the proposed framework. For different LLMs, we report the best training objective values achieved and their corresponding testing objective values.
Dataset GPT-4o (Baseline) GPT-4.5 GPT-4o-mini GPT-3.5
Train Test Train Test Train Test Train Test
Images -0.63 -0.48 -0.62 -0.54 -0.61 -0.48 -0.64 -0.52
Videos -1.63 -1.70 -1.62 -1.65 -1.62 -1.69 -1.63 -1.70
Time-series -0.42 -0.41 -0.40 -0.41 -0.41 -0.40 -0.42 -0.40
5 Conclusion
We have introduced tnLLM, a domain-aware LLM-guided TN-SS framework for directly infering TN structures. This has been achieved by utilizing the rich domain information in real-world tensor data and the inherent reasoning capabilities of LLMs. Experimental results have demonstrated that tnLLM achieves performance comparable to current SOTA algorithms, while requiring significantly fewer function evaluations. Notably, by incorporating domain information, our framework is the first to mitigate the black-box nature of the identified TN structures in TN-SS through generating domain-relevant solution explanations. Furthermore, we have shown that tnLLM can be used to accelerate SOTA sampling-based algorithms while preserving their theoretical guarantees.
Appendix A Limitations and future work
Despite the good performance, the proposed domain-aware TN-SS framework currently lacks theoretical guarantees for its evaluation efficiency due to the black-box nature of LLMs. This could be mitigated by developing domain-aware TN-SS algorithms based on continuous optimization. Also, as mentioned in the ablation study, weaker LLM models, such as GPT-3.5, were found to be more prone to misinterpret the optimization objectives. Therefore, improving the reasoning consistency of smaller LLM models promises to further improve the efficiency of the framework.
Appendix B Full prompts of tnLLM framework
The full “Behavior-directive”, “Task-directive” and “Optimization-directive” prompts used in the tnLLM framework are illustrated in Figures 8, 9, 10, respectively.
Figure 8: The full “Behavior-directive” prompt.
Figure 9: The full “Task-directive” prompt.
Figure 10: The full “Optimization-directive” prompt.
Appendix C Data
We constructed a custom financial time series dataset with 142 temporally ordered fifth-order tensors, denoted as $\{\mathcal{X}_{n}\}_{n=1}^{142}\in\mathbb{R}^{3\times 6\times 3\times 4\times 5}$, each representing a rolling window produced via multi-way delay embedding through the temporal direction [38]. This leads to the value selection process of $10$ ranks. To the best of our knowledge, this is the largest tensor dataset in terms of number of samples ever considered in the TN-SS problem. The first $80\%$ of these tensors were used as the training data, while the remaining $20\%$ with non-overlapping entries with the training data were used for testing. The modes of each time series tensor correspond to:
  • Mode 1: Types of financial instruments. They are equity indices, commodities, and currency swaps.
  • Mode 2: Assets within each type of financial instrument. For equity indices, these are Hang Seng, Nikkei 225, S&P 500, EURO STOXX 50, FTSE 100, and Shanghai Composite Index. For commodities, these are Brent Crude, Copper, Natural gas, Comex gold, Soybeans, and Wheat. For currency swaps, these are HKD/USD, JPY/USD, CHF/USD, EUR/USD, CNY/USD, and GBP/USD.
  • Mode 3: Features of each asset. These are average adjusted closing price log return, average relative price min-max, and average high-low spread.
  • Mode 4: Interval of time points on which we calculate the average features. There are 4 intervals – 1 day, 5 days, 10 days, and 15 days.
  • Mode 5: Time points within each rolling window of length 5.
The images and videos datasets are fetched from http://trace.eas.asu.edu/yuv/. Figure 11 illustrates 5 example image samples, $\{\mathcal{X}_{n}\}_{n=1}^{5}\in\mathbb{R}^{144\times 176\times 3}$, used in the experiments. The first $80\%$ of these tensors were used as the training data, while the remaining $20\%$ for testing. The modes of each image tensor correspond to:
  • Mode 1: Height of the image. These are the vertical pixel indexes.
  • Mode 2: Width of the image. These are the horizontal pixel indexes.
  • Mode 3: RGB channels. These are the red, green, and blue channels.
Figure 11: Image samples used in the experiments.
Video dataset is produced from performing the multi-way delay embedding [6] through the temporal mode in a video to produce 5 samples, $\{\mathcal{X}_{n}\}_{n=1}^{5}\in\mathbb{R}^{144\times 176\times 3\times 10}$. Figure 12 shows an example sample of the videos dataset at different frames. The modes of each video tensor correspond to:
  • Mode 1: Height of the image. These are the vertical pixel indexes.
  • Mode 2: Width of the image. These are the horizontal pixel indexes.
  • Mode 3: RGB channels. These are the red, green, and blue channels.
  • Mode 4: Frames in a video. These are the index of the frames in a video.
Figure 12: One example sample of the Video data used in the experiments.
Appendix D Implementation details
All baseline models are initialized as a “fully-disconnected” graph. The TN search template is set to a complete graph. For TNLS, we set the number of samples (evaluations), #$\#Sample$, in each sampling step to $4$ in the images dataset, $60$ in the videos dataset, and $10$ in the time series dataset. For TnALE, we set $D=1$, $r=1$. For tnGPS, we selected the best-performing algorithm discovered (“Ho-2”) and adopted the following hyperparameters as suggested in their paper: code upper bound of 10, mutation rate of 0.1, crossover rate of 0.6, selection pressure of 1.5, elitism enabled, diversity factor of 0.05, variance decay of 0.98, minimum variance of 0.1, tournament size factor of 0.2, elite diversity boost of 2.0, random individual chance of 0.05, and a maximum mutation of 3. All models were run for $500$ evaluations in both the images and videos dataset, and $250$ evaluations in the time series dataset if not converged. An internal server with NVIDIA RTX A6000 GPU, an AMD Ryzen Threadripper PRO 5955WX with 16 cores, and 256GB of RAM was used.
Appendix E Summarized generated explanations of tnLLM for the initialized TN structure
The full set of summarized LLM generated explanations over three independent runs for the images, videos, and time-series datasets are provided in Tables 3, 4, and 5, respectively. The suggested ranks are verified to be consistent with the domain-informed reasoning provided by tnLLM. Also, the domain-relevant explanations are robust across multiple runs.
Table 3: Summarized generated explanations of tnLLM for the initialized TN structure in the Images dataset over 3 independent runs.
Mode Pair Run 1 Run 2 Run 3
Mode 1 & Mode 2
(Width $\leftrightarrow$ Height)
Defines spatial resolution Spatial coherence is significant due to smoothness High rank needed to capture spatial detail (suggested rank: 20) Spatial dimensions are highly correlated Captures patterns and structures High rank to capture spatial complexity (suggested rank: 20) Spatial resolution key for details and patterns High rank crucial to capture more spatial features to maintain image quality High rank (suggested rank: 20)
Mode 1 & Mode 3
(Width $\leftrightarrow$ RGB)
Captures horizontal color distribution Less complex than spatial features Moderate rank to capture color variation without overfitting (suggested rank: 5) Horizontal color changes are smooth Color variations less complex as object maintains consistent color across width Moderate rank (suggested rank: 5) Related through color distribution along width Color variations less complex than spatial variation Moderate rank (suggested rank: 5)
Mode 2 & Mode 3
(Height $\leftrightarrow$ RGB)
Captures vertical color patterns Vertical and horizontal color distributions often equally important Rank should be similar to the width–RGB connection (suggested rank: 5) Captures vertical color variation Vertical structures (e.g., stripes) may add complexity Rank might be slightly higher than width–channel rank (suggested rank: 8) Complexity of color variation along height and width typically similar Moderate rank is also appropriate (suggested rank: 5)
Table 4: Summarized generated explanations of tnLLM for the initialized TN structure in the Videos dataset over 3 independent runs.
Mode Pair Run 1 Run 2 Run 3
Mode 1 & Mode 2
(Width $\leftrightarrow$ Height)
Defines spatial resolution per frame Needs to capture spatial detail for video quality but balance with compression Moderate-to-high rank (suggested rank: 20) Define spatial resolution per frame Higher rank captures more complex spatial patterns and correlations Moderate-to-high rank to balance detail capture and compression (suggested rank: 20) Define spatial resolution per frame Spatial resolution is a significant aspect of video quality Moderate-to-high rank to capture more spatial details across frames (suggested rank: 20)
Mode 1 & Mode 3, Mode 2 & Mode 3
(Width/Height $\leftrightarrow$ RGB)
Less complex than spatial info Limited variation of color across width/height Low rank suffices (suggested rank: 5) Less directly correlated than spatial dimensions Certain horizontal/vertical patterns prominent in specific color channels Low rank suffices (suggested rank: 5) Captures color variations horizontally/vertically Color variations can be significant in videos with rich color content Moderate-to-low rank (suggested rank: 5)
Mode 1 & Mode 4, Mode 2 & Mode 4
(Width/Height $\leftrightarrow$ Frames)
Captures horizontal/vertical structure changes over time Videos have temporal consistency Moderate rank (suggested rank: 10) Related through motion and changes across frames Moderate rank to capture temporal changes across width/height of the video (suggested rank: 10) Captures temporal variations across the horizontal/vertical dimension Temporal changes crucial for motion representation Moderate rank (suggested rank: 10)
Mode 3 & Mode 4
(RGB $\leftrightarrow$ Frames)
Captures color changes over time Color changes often subtle and less complex than spatial changes Lower rank can be used to maintain compression (suggested rank: 3) Captures color changes over time Crucial for capturing dynamic color variations and transitions Low rank to model temporal color variation (suggested rank: 5) Captures temporal color changes Essential for representing dynamic scenes Moderate-to-low rank necessary to capture these changes (suggested rank: 5)
Table 5: Summarized generated explanations of tnLLM for the initialized TN structure in the Time-series dataset over 3 independent runs.
Mode Pair Run 1 Run 2 Run 3
Mode 1 & Mode 2
(Types of financial instruments $\leftrightarrow$ Assets within each type)
Strong intrinsic relationship Different assets within a type exhibit correlated behaviors due to shared market factors Higher rank (suggested rank: 3) Strong relationship as each type is defined by its constituent assets High diversity and specificity across assets Higher rank (suggested rank: 3) Strong connection as each instrument has specific assets Rank should reflect the diversity and specificity of assets within each type Higher rank (suggested rank: 3)
Mode 1 & Mode 3
(Types of financial instruments $\leftrightarrow$ Features of each asset)
Features influenced by instrument type For example, volatility of commodities might differ from that of equity indexes Moderate rank (suggested rank: 2) Different instruments behave distinctly (returns, volatility, spreads) Features somewhat standardized across types Moderate rank to not overfit (suggested rank: 2) Instruments behave differently in terms of features However, features themselves are common across all types Moderate rank to capture variability (suggested rank: 2)
Mode 2 & Mode 3
(Assets within each type $\leftrightarrow$ Features of each asset)
Features directly derived from assets indicating strong relationship Different assets have distinct feature profiles Higher rank (suggested rank: 3) Each asset has unique characteristics in terms of returns, volatility, and spreads Higher rank to accurately capture specific features (suggested rank: 3) Each asset has unique characteristics in terms of features like log returns and spreads Higher rank needed for this detailed interaction (suggested rank: 3)
Mode 4 & Mode 5
(Interval of time points $\leftrightarrow$ Time points in each rolling window)
Inherently linked as rolling window defines the context for the time points Moderate rank (suggested rank: 2) Weak relationship as rolling window smooths out short-term variations Lower rank is suitable (suggested rank: 1) Intrinsic connection as time points are nested within the intervals Moderate rank to capture this structure (suggested rank: 2)
Appendix F Performance comparison for vanilla and tnLLM-initialized TNLS and TnALE
Table 6 illustrates the detailed performance of vanilla and tnLLM-initialized TNLS and TnALE across different datasets.
Table 6: Performance comparison across different datasets for vanilla and tnLLM-initialized TNLS and TnALE. The values on the left give the lowest training objective function values and their corresponding testing objective function values. The values in [square brackets] give the number of evaluations to first achieve the best training objective function value. For both metrics, a lower value is better. The least number of evaluations are highlighted in bold.
Data Type TNLS TnALE
Vanilla tnLLM-init Vanilla tnLLM-init
Images Train -0.66 [114] -0.64 [20] -0.65 [81] -0.64 [19]
Test -0.47 -0.46 -0.46 -0.46
Videos Train -1.64 [484] -1.64 [21] -1.66 [254] -1.65 [20]
Test -1.72 -1.71 -1.72 -1.70
Time-series Train -0.45 [218] -0.47 [14] -0.47 [177] -0.47 [20]
Test -0.43 -0.44 -0.47 -0.44
Appendix G Broader societal impact
This work uses LLMs to directly infer tensor network (TN) structures without human intervention by incorporating domain information about the data and utilizing the inherent reasoning capabilities of LLMs. The proposed framework significantly enhances efficiency and reduces computational costs compared to traditional methods, including sampling-based approaches which to date have demonstrated superior performance. By lowering computational demands, our approach also reduces $\text{CO}_{2}$ emissions and the environmental impact, a growing concern with big-data applications that often involve tensors. Furthermore, our approach provides domain-relevant explanations for the identified TN structures, enabling domain experts to verify the identified TN structures. This is particularly crucial in sensitive domains such as finance and healthcare.